# MaLA-500: Massive Language Adaptation of Large Language Models

### v1

@@ Appendix B Detailed Results
lim_Latn 49.49 43.43 49.49 50.51 25.25 47.47 60.61 61.62 60.61 66.67 66.67 64.65 71.72 44.44
lin_Latn 43.43 34.34 41.41 46.46 27.27 43.43 46.46 54.55 56.57 57.58 60.61 54.55 55.56 71.72
lit_Latn 38.38 46.46 39.39 45.45 19.19 50.51 58.59 62.63 65.66 64.65 66.67 63.64 65.66 53.54
lmo_Latn 52.53 48.48 53.54 53.54 19.19 54.55 60.61 54.55 60.61 60.61 60.61 60.61 65.66 63.64
ltz_Latn 56.57 48.48 50.51 50.51 32.32 53.54 65.66 63.64 68.69 70.71 70.71 65.66 72.73 62.63
lua_Latn 35.35 27.27 35.35 39.39 19.19 35.35 41.41 51.52 48.48 47.47 46.46 47.47 48.48 41.41
lug_Latn 28.28 27.27 38.38 36.36 18.18 31.31 41.41 40.40 42.42 44.44 44.44 39.39 44.44 46.46
luo_Latn 31.31 27.27 31.31 36.36 17.17 36.36 32.32 41.41 35.35 37.37 37.37 37.37 44.44 72.73
lus_Latn 42.42 36.36 43.43 44.44 20.20 40.40 43.43 47.47 43.43 47.47 47.47 44.44 49.49 66.67
lvs_Latn 47.47 57.58 44.44 43.43 30.30 59.60 67.68 66.67 64.65 69.70 72.73 68.69 76.77 69.70
mai_Deva 47.47 48.48 62.63 72.73 27.27 51.52 54.55 60.61 54.55 62.63 60.61 54.55 60.61 44.44
mal_Mlym 19.19 39.39 63.64 25.25 17.17 40.40 52.53 51.52 52.53 52.53 49.49 52.53 57.58 56.57
mar_Deva 34.34 47.47 65.66 48.48 21.21 46.46 56.57 60.61 57.58 55.56 60.61 55.56 60.61 72.73
min_Latn 51.52 48.48 59.60 55.56 21.21 52.53 59.60 66.67 56.57 67.68 65.66 58.59 65.66 60.61
mkd_Cyrl 56.57 44.44 39.39 55.56 46.46 57.58 61.62 65.66 68.69 71.72 74.75 70.71 72.73 71.72
mlt_Latn 46.46 43.43 45.45 43.43 20.20 58.59 70.71 68.69 69.70 72.73 74.75 74.75 78.79 68.69
mon_Cyrl 25.25 43.43 28.28 40.40 30.30 48.48 55.56 59.60 61.62 65.66 61.62 60.61 64.65 50.51
mos_Latn 24.24 26.26 30.30 35.35 18.18 29.29 36.36 38.38 35.35 39.39 32.32 37.37 40.4 52.53
mri_Latn 35.35 20.20 27.27 28.28 27.27 39.39 41.41 44.44 40.40 48.48 48.48 43.43 48.48 58.59
mya_Mymr 19.19 49.49 18.18 67.68 18.18 37.37 43.43 41.41 40.40 48.48 48.48 48.48 48.48 43.43
nld_Latn 53.54 47.47 53.54 64.65 38.38 60.61 70.71 74.75 69.70 73.74 77.78 72.73 75.76 75.76
nno_Latn 62.63 49.49 55.56 58.59 26.26 50.51 68.69 62.63 66.67 68.69 69.70 64.65 71.72 66.67
npi_Deva 50.51 41.41 62.63 61.62 28.28 51.52 60.61 65.66 62.63 69.70 66.67 63.64 66.67 67.68
nso_Latn 28.28 29.29 32.32 34.34 18.18 37.37 40.40 42.42 39.39 46.46 46.46 43.43 48.48 60.61
nya_Latn 34.34 33.33 40.40 43.43 22.22 37.37 44.44 53.54 54.55 56.57 56.57 56.57 50.51 50.51
oci_Latn 59.60 49.49 61.62 62.63 26.26 49.49 61.62 61.62 60.61 66.67 65.66 66.67 65.66 66.67
orm_Latn 22.22 23.23 22.22 23.23 14.14 26.26 29.29 35.35 37.37 32.32 38.38 37.37 38.38 58.59
ory_Orya 21.21 11.11 60.61 19.19 19.19 37.37 46.46 49.49 52.53 52.53 57.58 51.52 55.56 54.55
pag_Latn 57.58 47.47 55.56 56.57 29.29 57.58 62.63 69.70 66.67 67.68 67.68 69.7 68.69 57.58
pan_Guru 19.19 11.11 61.62 22.22 21.21 39.39 48.48 54.55 51.52 55.56 55.56 59.6 50.51 73.74
pap_Latn 57.58 49.49 54.55 58.59 25.25 59.60 70.71 70.71 69.70 69.70 72.73 73.74 73.74 71.72
pes_Arab 53.54 44.44 54.55 51.52 30.30 46.46 55.56 58.59 57.58 64.65 64.65 64.65 67.68 68.69
plt_Latn 30.30 30.30 33.33 31.31 20.20 45.45 50.51 52.53 54.55 56.57 60.61 58.59 65.66 74.75
pol_Latn 61.62 48.48 51.52 53.54 36.36 59.60 69.70 66.67 68.69 72.73 71.72 65.66 68.69 71.72
por_Latn 58.59 48.48 58.59 69.70 30.30 58.59 66.67 63.64 66.67 69.70 73.74 67.68 77.78 75.76
prs_Arab 54.55 34.34 48.48 51.52 27.27 49.49 55.56 60.61 60.61 64.65 57.58 61.62 64.65 68.69
pus_Arab 39.39 32.32 41.41 48.48 19.19 38.38 43.43 48.48 48.48 44.44 47.47 47.47 50.51 72.73
quy_Latn 40.40 37.37 43.43 46.46 23.23 43.43 46.46 54.55 47.47 51.52 52.53 48.48 52.53 74.75
ron_Latn 64.65 53.54 60.61 57.58 32.32 57.58 63.64 64.65 67.68 71.72 69.70 63.64 72.73 71.72
run_Latn 29.29 30.30 43.43 33.33 17.17 42.42 52.53 56.57 49.49 55.56 51.52 54.55 54.55 66.67
rus_Cyrl 62.63 51.52 53.54 63.64 43.43 60.61 68.69 67.68 67.68 70.71 73.74 68.69 75.76 74.75
sag_Latn 26.26 29.29 34.34 40.40 24.24 37.37 39.39 44.44 44.44 47.47 43.43 45.45 49.49 44.44
san_Deva 38.38 45.45 52.53 51.52 22.22 41.41 53.54 49.49 42.42 48.48 49.49 42.42 39.39 68.69
scn_Latn 43.43 45.45 52.53 51.52 26.26 60.61 67.68 66.67 66.67 68.69 74.75 70.71 75.76 36.36
sin_Sinh 22.22 09.09 12.12 16.16 23.23 49.49 51.52 53.54 55.56 60.61 59.60 60.61 57.58 63.64
slk_Latn 63.64 52.53 48.48 49.49 32.32 56.57 62.63 62.63 66.67 69.70 68.69 67.68 68.69 63.64
slv_Latn 65.66 53.54 42.42 49.49 32.32 57.58 64.65 65.66 67.68 68.69 70.71 70.71 74.75 63.64
smo_Latn 30.30 30.30 31.31 35.35 24.24 49.49 59.60 60.61 59.60 61.62 65.66 59.6 65.66 76.77
sna_Latn 32.32 26.26 41.41 35.35 21.21 36.36 41.41 43.43 42.42 46.46 47.47 39.39 40.4 68.69
snd_Arab 38.38 35.35 34.34 43.43 25.25 42.42 37.37 45.45 40.40 49.49 47.47 45.45 47.47 55.56
som_Latn 32.32 29.29 26.26 38.38 15.15 31.31 41.41 41.41 47.47 44.44 42.42 45.45 46.46 66.67
sot_Latn 29.29 24.24 35.35 33.33 19.19 40.40 44.44 49.49 50.51 49.49 52.53 51.52 54.55 73.74
spa_Latn 59.60 51.52 66.67 67.68 33.33 58.59 69.70 69.70 67.68 72.73 71.72 73.74 76.77 60.61
srd_Latn 45.45 40.40 49.49 54.55 19.19 54.55 65.66 66.67 62.63 68.69 73.74 66.67 72.73 51.52
srp_Cyrl 62.63 48.48 45.45 59.60 49.49 58.59 64.65 58.59 67.68 69.70 69.70 63.64 69.7 66.67
ssw_Latn 25.25 27.27 33.33 29.29 18.18 36.36 39.39 51.52 49.49 48.48 52.53 47.47 48.48 61.62
sun_Latn 53.54 49.49 57.58 60.61 31.31 59.60 68.69 68.69 67.68 67.68 68.69 66.67 73.74 66.67
swe_Latn 65.66 49.49 52.53 61.62 29.29 57.58 66.67 64.65 66.67 72.73 72.73 73.74 75.76 62.63
swh_Latn 43.43 51.52 60.61 70.71 29.29 53.54 61.62 61.62 58.59 62.63 56.57 62.63 66.67 68.69
szl_Latn 51.52 43.43 44.44 41.41 35.35 52.53 56.57 59.60 64.65 63.64 68.69 62.63 71.72 54.55
tam_Taml 29.29 47.47 59.60 64.65 21.21 39.39 47.47 51.52 55.56 65.66 58.59 57.58 60.61 74.75
tat_Cyrl 36.36 56.57 34.34 47.47 30.30 50.51 55.56 58.59 59.60 61.62 63.64 63.64 67.68 68.69
[…]

### v2

@@ Appendix B Detailed Results
lim_Latn 60.78 50.00 50.00 33.82 32.84 56.37 60.78 63.24 66.67 67.16 69.12 72.55 70.59 72.06
lin_Latn 36.76 40.20 43.14 34.31 23.04 38.24 47.06 53.92 57.84 61.27 56.86 60.29 62.75 65.69
lit_Latn 40.20 60.29 41.18 30.39 32.35 55.39 62.75 65.20 64.71 70.59 68.14 66.67 69.61 73.04
lmo_Latn 57.84 50.98 55.88 41.67 34.80 59.31 65.69 66.18 70.10 71.57 70.59 70.59 75.00 75.49
ltz_Latn 55.88 47.06 52.94 39.22 39.22 56.37 65.20 61.27 70.59 68.14 71.08 70.59 71.08 74.02
lua_Latn 32.35 33.33 39.22 28.43 20.10 33.82 40.20 42.65 49.02 51.96 50.00 50.00 49.51 50.00
lug_Latn 27.94 25.00 33.82 19.61 22.06 35.78 40.20 43.63 48.04 51.47 47.06 43.63 49.02 49.02
luo_Latn 28.43 28.43 32.84 25.49 21.57 31.37 37.25 42.65 48.04 49.51 46.57 51.47 49.51 51.47
lus_Latn 43.63 42.16 49.02 31.37 25.49 45.59 51.96 53.92 52.45 54.90 57.84 60.29 58.33 60.29
lvs_Latn 43.14 67.16 43.63 29.41 31.37 57.84 65.20 63.24 68.14 72.55 67.65 69.61 71.08 72.55
mai_Deva 40.69 59.31 60.29 51.47 33.33 57.84 61.76 66.67 67.65 69.12 69.12 70.10 71.57 69.12
mal_Mlym 20.10 60.29 64.71 13.24 25.98 52.45 59.31 60.29 62.75 62.25 65.69 63.24 63.73 68.14
mar_Deva 29.90 56.86 63.73 37.75 36.27 51.96 57.35 64.22 63.73 63.73 63.73 66.67 66.18 68.14
min_Latn 48.04 55.39 59.80 39.71 31.37 57.35 69.12 68.14 68.63 77.94 72.06 75.98 75.49 77.45
mkd_Cyrl 60.78 52.45 32.84 44.12 44.12 66.18 68.63 69.12 68.63 73.04 73.04 72.55 73.04 76.96
mlt_Latn 49.51 45.10 46.08 29.90 35.78 64.71 67.16 68.14 67.16 77.45 75.49 76.96 77.45 76.96
mon_Cyrl 23.53 54.90 20.10 18.63 38.24 50.00 56.86 55.88 63.24 64.22 63.24 63.24 65.20 67.16
mos_Latn 25.49 23.53 29.90 20.59 20.59 27.94 36.76 37.75 37.75 40.69 41.67 45.10 41.67 45.10
mri_Latn 30.39 24.02 30.88 17.65 28.43 44.12 49.02 51.47 51.47 57.84 55.88 58.82 56.37 58.33
mya_Mymr 19.12 60.29 19.61 60.29 23.53 38.73 43.14 53.43 53.43 50.98 52.45 54.90 51.96 54.90
nld_Latn 70.10 59.80 55.88 46.08 45.59 64.71 69.12 68.63 73.04 73.53 75.49 74.02 79.41 80.88
nno_Latn 64.71 61.76 52.45 45.59 35.29 52.94 64.22 62.75 66.18 68.63 68.63 70.10 69.12 73.04
npi_Deva 39.22 51.96 64.71 40.69 33.82 57.84 61.76 68.14 67.65 67.65 68.63 70.10 68.63 75.49
nso_Latn 27.94 30.88 33.33 22.55 21.08 33.82 43.14 46.08 49.02 52.94 51.96 53.92 54.90 53.92
nya_Latn 32.35 34.31 40.69 27.94 23.04 35.29 45.59 49.02 50.98 51.47 52.94 53.92 52.94 58.33
oci_Latn 68.63 56.37 65.69 48.53 34.31 60.29 69.12 65.20 67.65 73.04 73.53 71.57 75.49 76.47
orm_Latn 17.16 18.14 22.06 16.67 20.10 30.39 35.29 41.18 41.67 47.55 41.67 44.12 43.14 51.47
ory_Orya 13.24 13.73 64.22 11.76 24.51 45.10 52.45 57.84 53.92 61.27 57.84 56.86 60.78 60.78
pag_Latn 52.45 49.51 53.92 40.20 31.86 54.90 62.75 60.78 67.65 64.71 70.10 70.10 69.12 69.61
pan_Guru 14.22 11.27 62.25 11.76 33.82 54.90 58.82 63.73 64.22 67.65 67.16 66.67 68.63 67.16
pap_Latn 55.39 50.00 52.94 38.24 30.39 56.86 64.71 66.67 69.61 74.51 69.12 73.53 70.59 75.49
pes_Arab 47.06 58.82 52.94 32.84 39.22 61.27 71.08 63.73 70.59 72.55 72.55 73.53 76.47 76.47
plt_Latn 28.43 32.84 37.25 21.57 29.41 51.96 58.82 57.84 59.31 60.78 60.29 60.29 64.22 60.29
pol_Latn 74.51 60.78 47.06 32.84 36.76 61.76 68.63 69.12 71.08 75.00 74.02 74.02 77.45 75.98
por_Latn 70.10 61.76 65.20 59.31 36.76 64.71 72.06 70.10 74.51 75.00 76.96 75.49 78.43 82.84
prs_Arab 50.49 55.39 49.51 33.33 37.25 60.78 64.22 67.16 69.12 72.55 72.55 73.53 72.55 75.49
pus_Arab 30.39 34.80 38.73 21.08 30.39 47.06 50.98 52.45 54.41 53.92 53.92 55.88 55.88 57.84
quy_Latn 32.84 35.29 40.69 35.29 22.06 36.27 44.12 45.59 49.02 52.45 49.51 49.02 50.98 50.98
ron_Latn 69.12 61.76 57.84 42.65 41.18 61.27 70.10 65.20 70.10 74.51 73.53 75.00 78.92 78.43
run_Latn 25.49 27.94 44.12 25.49 23.53 37.25 46.57 50.49 51.96 59.31 51.96 56.37 57.84 60.29
rus_Cyrl 71.57 63.24 53.43 60.29 38.73 64.22 65.20 69.12 72.06 75.98 75.00 76.47 75.49 78.92
sag_Latn 29.90 27.94 31.37 21.08 20.59 30.88 43.63 47.06 48.53 55.88 52.45 54.41 55.88 58.82
san_Deva 27.94 47.55 54.90 42.65 24.51 48.04 60.29 57.84 62.25 66.67 65.20 61.76 66.18 65.20
scn_Latn 51.96 50.00 53.43 40.69 37.25 63.73 73.04 70.59 74.02 77.45 75.49 75.00 80.39 76.47
sin_Sinh 15.20 10.78 20.10 12.75 29.90 56.37 60.29 65.20 66.18 68.14 64.71 66.67 63.73 67.16
slk_Latn 68.14 60.29 47.55 39.71 34.31 58.33 68.63 66.67 70.59 75.00 70.59 71.57 74.51 75.00
slv_Latn 68.14 60.78 44.12 32.84 38.73 63.24 68.14 68.14 70.59 73.53 73.53 74.51 78.43 76.47
smo_Latn 30.39 25.00 31.86 18.14 29.41 52.45 60.29 62.25 62.25 65.69 67.16 65.20 66.18 69.61
sna_Latn 28.43 29.41 36.27 23.53 24.51 39.71 44.61 45.59 44.61 49.51 45.59 47.55 47.06 50.00
snd_Arab 27.94 37.25 39.22 23.53 27.94 42.65 47.06 50.49 52.45 54.41 54.41 52.94 55.88 56.86
som_Latn 23.53 25.49 27.94 17.16 22.06 36.27 44.61 47.55 51.47 52.94 52.94 53.92 54.41 55.39
sot_Latn 29.41 28.43 33.82 18.63 22.55 36.76 43.14 47.06 50.49 51.96 52.45 55.39 54.41 56.86
spa_Latn 72.55 58.33 67.65 56.37 35.29 64.22 69.61 72.06 74.51 74.02 72.06 76.47 78.43 78.43
srd_Latn 53.92 52.45 50.98 37.25 31.37 60.29 66.18 68.63 75.98 74.02 77.94 77.45 79.41 79.41
srp_Cyrl 63.73 55.39 33.33 39.22 45.59 65.20 70.59 69.61 73.04 76.47 74.02 75.00 77.94 79.41
ssw_Latn 29.41 25.00 31.37 21.57 24.02 44.12 46.57 50.00 52.94 51.96 53.92 56.86 53.92 60.78
sun_Latn 55.39 59.31 63.73 44.61 37.25 60.29 68.63 70.10 71.08 73.53 72.55 73.53 75.00 75.98
swe_Latn 71.08 61.27 52.94 48.04 33.82 53.43 60.29 64.71 64.71 69.12 70.10 69.61 72.55 70.59
swh_Latn 32.35 63.24 61.27 56.86 29.41 50.49 59.31 58.82 62.75 60.29 62.25 68.63 66.67 66.67
szl_Latn 56.86 50.49 45.59 29.41 30.88 51.47 59.80 63.73 64.71 67.16 69.61 68.63 71.08 69.12
tam_Taml 20.59 63.24 67.16 58.82 30.88 50.49 55.88 62.75 63.24 63.73 65.20 69.61 66.67 68.63
tat_Cyrl 37.75 60.29 35.29 28.92 33.33 54.90 64.22 64.71 65.69 74.51 70.10 71.57 71.57 73.53
tel_Telu 18.14 60.78 61.27 59.80 25.98 50.00 52.45 59.80 58.82 63.73 60.78 60.29 63.73 61.76
tgk_Cyrl 26.96 57.84 23.53 17.16 36.76 54.90 60.29 60.78 61.27 69.12 64.71 66.18 68.14 70.59
tgl_Latn 55.88 58.33 49.02 40.20 43.14 64.22 69.61 64.71 70.10 75.49 74.51 77.45 78.43 77.45
tha_Thai 44.61 60.78 23.53 57.35 41.18 63.24 67.16 68.63 70.10 72.06 70.59 70.10 72.06 75.49
tir_Ethi 13.24 16.18 16.18 13.73 21.57 34.80 39.22 41.67 47.06 46.08 45.59 47.06 45.59 47.55
tpi_Latn 63.24 46.57 56.86 33.33 31.86 58.82 65.69 68.14 70.10 72.06 74.51 73.53 75.98 74.51
tsn_Latn 28.92 29.90 32.35 24.51 24.51 40.20 44.12 47.06 47.06 53.43 51.96 50.00 50.49 54.41
tso_Latn 30.88 31.37 36.76 28.92 22.55 35.29 41.18 45.10 46.08 48.04 43.14 43.14 45.59 49.02
tuk_Latn 34.31 46.08 39.22 27.45 24.02 45.59 53.43 58.82 57.84 63.73 63.73 66.18 65.69 66.18
tum_Latn 26.96 34.31 33.82 27.94 21.57 39.22 43.14 45.59 46.08 47.55 44.61 49.02 49.51 49.51
tur_Latn 52.94 62.75 40.20 52.94 36.76 60.78 68.63 70.10 72.06 74.02 75.00 76.47 75.98 76.96
uig_Arab 18.63 18.14 20.10 11.27 21.08 33.33 36.76 39.71 43.14 44.61 43.14 48.53 47.55 48.04
ukr_Cyrl 71.57 63.73 41.18 43.63 39.71 60.29 65.69 66.18 69.12 71.08 75.00 73.53 72.55 75.00
umb_Latn 25.00 26.47 29.90 23.04 21.57 30.88 32.84 36.76 35.78 40.20 38.24 34.80 36.76 35.29
urd_Arab 38.73 53.43 63.24 54.41 36.27 55.39 62.75 64.22 64.22 68.63 65.20 68.63 67.16 67.16
uzb_Latn 30.39 62.75 35.78 23.53 22.06 50.49 56.37 57.84 63.24 72.06 63.73 69.12 72.55 71.57
vec_Latn 65.69 59.80 56.86 52.45 39.22 62.25 66.18 69.61 70.10 69.12 74.51 75.98 75.00 76.47
vie_Latn 67.65 63.24 67.16 60.78 39.71 60.78 67.16 68.63 74.51 75.98 76.47 75.00 78.43 79.90
war_Latn 51.47 51.47 51.47 37.25 37.75 61.27 65.69 65.20 69.61 73.04 71.57 71.57 74.02 74.02
wol_Latn 32.35 34.80 43.14 25.49 23.53 36.76 42.16 45.59 48.53 53.43 47.55 52.94 54.90 53.43
xho_Latn 30.39 29.90 38.24 22.55 25.98 46.57 51.96 56.37 58.33 60.78 61.76 60.29 62.75 64.22
yid_Hebr 23.04 22.06 16.18 12.25 24.02 34.80 39.22 39.71 40.20 46.57 41.18 40.69 44.61 45.10
yor_Latn 21.57 29.41 47.55 21.57 26.47 32.35 41.18 39.22 41.67 48.04 42.16 43.14 44.61 43.14
yue_Hani 75.00 64.71 67.16 55.88 40.69 69.12 71.57 76.47 76.96 81.37 77.94 79.41 81.86 79.41
zsm_Latn 65.69 61.27 64.71 50.00 36.76 60.29 68.14 69.12 67.65 73.53 73.53 76.96 77.45 75.00
zul_Latn 25.00 25.49 35.29 15.20 25.49 51.47 49.02 54.41 56.37 57.84 60.29 61.76 59.31 62.75
all 42.08 45.34 44.63 34.36 30.88 50.71 57.02 58.95 61.20 64.04 63.15 64.13 65.19 66.32
Table 23: Detailed results on SIB200 (Part II). For previous LLMs, 3-shot results are presented.
Lang LLaMA 2-7B mGPT-13B BLOOM-7B1 XGLM-7.5B MaLA-500
ace_Latn 46.85 47.75 49.55 41.44 48.65
ach_Latn 45.05 37.84 41.44 40.54 36.04
acr_Latn 47.75 51.35 50.45 47.75 50.45
afr_Latn 54.05 38.74 51.35 49.55 55.86
agw_Latn 48.65 45.05 41.44 42.34 49.55
ahk_Latn 43.24 36.04 36.04 35.14 45.05
aka_Latn 42.34 32.43 38.74 42.34 54.95
aln_Latn 34.23 35.14 36.94 35.14 44.14
als_Latn 38.74 36.94 42.34 42.34 47.75
alt_Cyrl 44.14 44.14 45.05 51.35 51.35
alz_Latn 36.94 35.14 31.53 28.83 37.84
aoj_Latn 50.93 37.96 45.37 46.3 49.07
arb_Arab 43.24 45.05 49.55 44.14 50.45
arn_Latn 38.74 42.34 34.23 36.04 43.24
ary_Arab 32.43 33.33 38.74 32.43 44.14
arz_Arab 31.53 39.64 45.05 36.94 46.85
asm_Beng 45.95 42.34 54.95 40.54 54.05
ayr_Latn 47.75 37.84 44.14 44.14 54.05
azb_Arab 39.64 40.54 47.75 45.05 47.75
aze_Latn 45.05 46.85 45.05 43.24 49.55
bak_Cyrl 45.05 52.25 49.55 56.76 56.76
bam_Latn 42.34 37.84 49.55 39.64 47.75
ban_Latn 36.04 41.44 34.23 34.23 42.34
bar_Latn 49.55 46.85 44.14 48.65 53.15
bba_Latn 45.05 32.43 45.95 46.85 46.85
bci_Latn 36.94 35.14 36.94 33.33 44.14
bcl_Latn 42.34 48.65 39.64 39.64 54.95
bel_Cyrl 47.75 45.95 48.65 43.24 57.66
bem_Latn 47.75 37.84 42.34 41.44 51.35
ben_Beng 40.54 41.44 52.25 51.35 47.75
bhw_Latn 37.84 43.24 41.44 46.85 47.75
bim_Latn 38.74 39.64 33.33 36.94 45.05
bis_Latn 44.14 49.55 44.14 39.64 48.65
bqc_Latn 39.64 36.04 34.23 33.33 40.54
bre_Latn 39.64 36.04 35.14 36.04 40.54
btx_Latn 49.55 36.94 42.34 41.44 43.24
bul_Cyrl 45.05 42.34 48.65 45.05 54.95
bum_Latn 42.34 39.64 37.84 37.84 44.14
bzj_Latn 53.15 46.85 47.75 50.45 52.25
cab_Latn 39.64 38.74 37.84 36.94 36.04
cac_Latn 43.24 37.84 40.54 38.74 45.05
cak_Latn 45.95 35.14 44.14 40.54 50.45
caq_Latn 39.64 38.74 38.74 44.14 37.84
cat_Latn 52.25 45.05 46.85 48.65 52.25
cbk_Latn 54.05 40.54 56.76 54.05 55.86
cce_Latn 49.55 45.05 50.45 48.65 48.65
ceb_Latn 44.14 42.34 48.65 45.05 51.35
ces_Latn 44.14 43.24 45.05 46.85 51.35
cfm_Latn 49.55 41.44 49.55 53.15 48.65
che_Cyrl 37.84 33.33 36.94 37.84 38.74
chk_Latn 45.05 41.44 41.44 36.04 45.95
chv_Cyrl 43.24 45.05 45.05 49.55 58.56
ckb_Arab 44.14 36.94 45.05 42.34 51.35
cmn_Hani 48.65 45.05 53.15 48.65 53.15
cnh_Latn 46.85 46.85 46.85 49.55 46.85
crh_Cyrl 49.55 40.54 47.75 54.95 54.05
crs_Latn 52.25 44.14 49.55 55.86 59.46
csy_Latn 47.75 41.44 54.95 53.15 45.95
ctd_Latn 50.45 48.65 56.76 53.15 56.76
ctu_Latn 41.44 35.14 38.74 40.54 43.24
cuk_Latn 42.34 42.34 38.74 39.64 37.84
cym_Latn 39.64 38.74 39.64 43.24 41.44
dan_Latn 53.15 41.44 39.64 38.74 54.95
deu_Latn 45.05 36.04 37.84 38.74 43.24
djk_Latn 42.34 35.14 42.34 46.85 40.54
dln_Latn 48.65 40.54 51.35 54.05 47.75
Table 24: Detailed results on Taxi1500 (Part I). 3-shot results are presented.
Lang LLaMA 2-7B mGPT-13B BLOOM-7B1 XGLM-7.5B MaLA-500
dtp_Latn 39.64 35.14 42.34 46.85 50.45
dyu_Latn 41.44 39.64 42.34 38.74 46.85
dzo_Tibt 45.05 40.54 41.44 45.05 45.05
efi_Latn 39.64 36.04 38.74 41.44 45.95
ell_Grek 49.55 45.95 49.55 48.65 51.35
eng_Latn 55.86 42.34 58.56 54.05 59.46
enm_Latn 50.45 41.44 56.76 50.45 54.95
epo_Latn 49.55 40.54 47.75 42.34 49.55
est_Latn 46.85 42.34 38.74 52.25 46.85
eus_Latn 38.74 36.04 36.94 39.64 39.64
ewe_Latn 51.35 43.24 50.45 46.85 45.05
fao_Latn 53.15 44.14 52.25 53.15 58.56
fas_Arab 49.55 50.45 57.66 51.35 55.86
fij_Latn 48.65 43.24 41.44 43.24 53.15
fil_Latn 48.65 41.44 46.85 51.35 51.35
fin_Latn 47.75 45.05 41.44 45.95 54.95
fon_Latn 38.74 35.14 37.84 40.54 45.05
fra_Latn 60.36 51.35 62.16 52.25 59.46
fry_Latn 37.84 33.33 36.04 27.03 46.85
gaa_Latn 41.44 33.33 37.84 35.14 40.54
gil_Latn 36.7 31.19 41.28 32.11 41.28
giz_Latn 46.85 44.14 43.24 38.74 45.05
gkn_Latn 38.74 34.23 34.23 36.94 41.44
gkp_Latn 30.63 33.33 41.44 29.73 48.65
gla_Latn 33.33 39.64 44.14 45.05 49.55
gle_Latn 33.33 35.14 36.04 34.23 39.64
glv_Latn 43.24 41.44 37.84 38.74 42.34
gom_Latn 34.23 31.53 33.33 40.54 42.34
gor_Latn 43.24 34.23 43.24 40.54 46.85
guc_Latn 44.14 36.04 37.84 41.44 45.05
gug_Latn 45.05 44.14 42.34 41.44 50.45
guj_Gujr 45.95 37.84 52.25 44.14 56.76
gur_Latn 45.95 45.95 44.14 47.75 48.65
guw_Latn 45.05 37.84 47.75 46.85 48.65
gya_Latn 37.84 37.84 41.44 34.23 42.34
gym_Latn 41.44 39.64 39.64 43.24 50.45
hat_Latn 50.45 43.24 44.14 41.44 56.76
hau_Latn 44.14 37.84 41.44 44.14 48.65
haw_Latn 45.95 39.64 38.74 34.23 49.55
heb_Hebr 38.74 35.14 34.23 36.94 44.14
hif_Latn 42.34 43.24 49.55 47.75 48.65
hil_Latn 49.55 41.44 40.54 36.94 54.95
hin_Deva 51.35 50.45 49.55 46.85 56.76
hmo_Latn 46.85 45.05 46.85 45.05 53.15
hne_Deva 55.86 54.05 54.05 58.56 58.56
hnj_Latn 48.65 45.05 53.15 51.35 60.36
hra_Latn 49.55 41.44 43.24 46.85 45.95
hrv_Latn 55.86 51.35 52.25 54.95 61.26
hui_Latn 51.35 40.54 41.44 45.05 46.85
hun_Latn 46.85 44.14 41.44 43.24 48.65
hus_Latn 32.43 32.43 34.23 37.84 42.34
hye_Armn 45.95 40.54 45.95 49.55 61.26
iba_Latn 49.55 46.85 51.35 48.65 55.86
ibo_Latn 38.74 33.33 43.24 38.74 44.14
ifa_Latn 36.04 30.63 35.14 38.74 43.24
ifb_Latn 34.23 35.14 39.64 34.23 53.15
ikk_Latn 43.24 36.94 39.64 39.64 43.24
ilo_Latn 39.64 36.04 41.44 37.84 43.24
ind_Latn 49.55 50.45 53.15 53.15 54.95
isl_Latn 48.65 44.14 43.24 48.65 54.05
ita_Latn 50.45 49.55 56.76 58.56 54.95
ium_Latn 45.95 44.14 49.55 50.45 45.05
ixl_Latn 42.34 39.64 41.44 43.24 40.54
izz_Latn 38.74 47.75 39.64 42.34 54.95
jam_Latn 41.44 43.24 53.15 50.45 61.26
jav_Latn 41.44 47.75 44.14 37.84 45.95
Table 25: Detailed results on Taxi1500 (Part II). 3-shot results are presented.
Lang LLaMA 2-7B mGPT-13B BLOOM-7B1 XGLM-7.5B MaLA-500
jpn_Jpan 46.85 46.85 47.75 50.45 51.35
kaa_Latn 43.24 53.15 47.75 51.35 54.05
kab_Latn 27.93 36.04 30.63 34.23 35.14
kac_Latn 44.14 34.23 43.24 42.34 52.25
kal_Latn 41.44 37.84 36.04 35.14 40.54
kan_Knda 48.65 37.84 52.25 45.95 54.05
kat_Geor 41.44 41.44 42.34 46.85 48.65
kaz_Cyrl 49.55 45.05 51.35 53.15 55.86
kbp_Latn 40.54 35.14 36.94 31.53 47.75
kek_Latn 45.95 42.34 45.05 44.14 51.35
khm_Khmr 52.25 38.74 48.65 49.55 64.86
kia_Latn 36.94 36.04 40.54 41.44 48.65
kik_Latn 45.05 43.24 45.05 44.14 50.45
kin_Latn 42.34 37.84 41.44 38.74 50.45
kir_Cyrl 51.35 46.85 47.75 63.06 64.86
kjb_Latn 48.65 46.85 44.14 44.14 48.65
kjh_Cyrl 44.14 41.44 45.05 41.44 45.95
kmm_Latn 45.95 45.05 47.75 51.35 45.95
kmr_Cyrl 39.64 35.14 45.05 42.34 43.24
knv_Latn 44.55 44.55 45.45 42.73 44.55
kor_Hang 48.65 48.65 49.55 51.35 62.16
kpg_Latn 44.14 52.25 51.35 42.34 54.95
krc_Cyrl 45.95 36.04 48.65 48.65 53.15
kri_Latn 49.55 48.65 49.55 51.35 54.95
ksd_Latn 36.94 33.33 40.54 33.33 49.55
kss_Latn 32.43 28.83 34.23 29.73 47.75
ksw_Mymr 44.14 45.95 42.34 37.84 52.25
kua_Latn 41.44 42.34 36.94 35.14 40.54
lam_Latn 43.24 36.94 45.95 43.24 40.54
lao_Laoo 45.05 39.64 46.85 50.45 50.45
lat_Latn 53.15 41.44 53.15 56.76 57.66
lav_Latn 39.64 33.33 36.04 39.64 45.05
ldi_Latn 35.14 32.43 36.94 34.23 36.04
leh_Latn 47.75 37.84 33.33 32.43 41.44
lhu_Latn 27.93 34.23 34.23 37.84 42.34
lin_Latn 47.75 37.84 39.64 39.64 48.65
lit_Latn 42.34 40.54 44.14 48.65 49.55
loz_Latn 45.95 42.34 36.04 44.14 40.54
ltz_Latn 46.85 45.95 47.75 41.44 49.55
lug_Latn 40.54 32.43 39.64 38.74 45.95
luo_Latn 40.54 36.94 34.23 38.74 40.54
lus_Latn 39.64 40.54 42.34 41.44 50.45
lzh_Hani 54.95 48.65 54.05 43.24 56.76
mad_Latn 47.75 52.25 47.75 47.75 53.15
mah_Latn 43.24 36.04 42.34 45.95 45.05
mai_Deva 45.05 41.44 49.55 54.05 51.35
mam_Latn 43.24 33.33 41.44 45.05 45.95
mar_Deva 49.55 44.14 53.15 45.95 56.76
mau_Latn 29.73 29.73 36.94 37.84 32.43
mbb_Latn 44.14 42.34 38.74 39.64 49.55
mck_Latn 40.54 34.23 36.04 39.64 49.55
mcn_Latn 35.14 27.93 33.33 33.33 38.74
mco_Latn 41.44 33.33 43.24 33.33 43.24
[…]